OpenStudy (anonymous):

Give the complete factorization for the following polynomial. 75n + 3n^3

OpenStudy (anonymous):

is that \[\huge 75n+3n^3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

factor out 3n

OpenStudy (anonymous):

how do I facor with 3n^3??

OpenStudy (anonymous):

ask your self what happens when you divide the expression by 3n whats left???

OpenStudy (anonymous):

um n?

OpenStudy (anonymous):

3n( + )

OpenStudy (anonymous):

3n(?+?)=

OpenStudy (anonymous):

the problem asks for a format like this: ___n^__(n^__ + ___) if this helps us any

OpenStudy (anonymous):

yes we start by factoring 3n first then get to that form 1st you shud try to do this 1st

OpenStudy (anonymous):

3n(25n + n) is that right?

OpenStudy (anonymous):

@kljasdjasdl why are you giving me a picture that has to do with angles?

OpenStudy (anonymous):

3n(25+n^2) right dont forget n^2

OpenStudy (anonymous):

also 25 without n

OpenStudy (anonymous):

oh, now I get it. divide 3n^3 by 3n and you lose one 3n

OpenStudy (anonymous):

@kljasdjasdl use \[s=r\theta\] \[3/4\pi=4\theta\]

OpenStudy (anonymous):

yes @craftshark98 correct

OpenStudy (anonymous):

we know have \[\huge{3n(25+n^2)}\] here we can ongly here theres nothing more we can do than write complex factors

OpenStudy (anonymous):

are you sure the problem is not 75n-3n^3 but 75n+3n^3

OpenStudy (anonymous):

is it + or -

OpenStudy (anonymous):

I'm pretty sure it's an addition sign. I'm looking right at it.

OpenStudy (anonymous):